# Know When to Abstain: Optimal Selective Classification with Likelihood Ratios

### v1

@@ 1 Introduction
A wide range of methods have been proposed to determine whether a model should accept or reject an input. Common post-hoc approaches rely on heuristic confidence estimates, such as the maximum softmax probability [1, 2], logit margins [3], or Monte Carlo dropout [1, 4]. Other techniques assess a sample’s proximity to the training distribution [5, 6], under the assumption that samples farther from the data manifold are more likely to be misclassified. A separate line of work trains models with explicit abstention mechanisms, such as a rejection logit or head [7, 8, 9]. In this work we focus on post-hoc methods, which are model-agnostic and do not require specialized training.
Despite the rich literature, two important gaps remain. First, while foundational results (such as [10, 1]) provide theoretical underpinnings for selective classification, there is a lack of general, principled guidance for designing effective selector functions in the context of modern deep networks. Second, most evaluations are conducted in the i.i.d. setting, where test data is assumed to follow the training distribution. This setup fails to reflect challenges in real-world scenarios, where distribution shifts, especially covariate shifts, are common, reducing the practical relevance of existing results.
To address these challenges, we propose a new perspective rooted in the Neyman–Pearson lemma, a classical result from statistics that defines the optimal hypothesis test in terms of a likelihood ratio. We show that existing selectors can be interpreted as approximations to this test, and we use this insight to derive two new selectors, $\Delta$-MDS and $\Delta$-KNN, as well as a simple linear combination strategy. We evaluate our methods on a comprehensive suite of vision and language benchmarks under covariate shift, where the input distribution changes while the label space remains fixed. We focus on covariate shift for two key reasons: first, it is underexplored relative to semantic shifts [11] (which is well studied in the context of out-of-distribution detection [2, 12, 5, 13]); second, it is increasingly relevant in modern applications such as vision-language models (VLMs), where the label set is large and variable, rendering most practical shifts in deployment covariate in nature. Our results demonstrate that the proposed selectors outperform existing baselines and provide robust performance across distribution shifts, including on powerful VLMs like CLIP.
@@
    We conduct a thorough evaluation under covariate shift across vision and language tasks and demonstrate superior performance across both VLMs and traditional supervised models.
@@ 2 Background
where $s:\mathcal{X}\rightarrow\mathbb{R}$ is a confidence scoring function (often adapted from OOD detection methods; see below), and $\gamma$ is a tunable threshold.
The performance of a selective classifier is typically evaluated using two metrics:
@@
where $\ell(f({\bm{x}}),y)$ is the 0/1 loss [1]. Selective classification aims to optimize the tradeoff between selective risk and coverage by ideally reducing risk while maintaining high coverage. Improvements can stem from enhancing the base classifier $f$, or from refining the selector $g$ to better identify error-prone inputs. In this work, we fix $f$ to be a strong pretrained model and focus on designing more effective selector functions $g$.
@@ 3 Selective Classification via the Neyman–Pearson Lemma
We begin by framing selective classification within the classical paradigm of hypothesis testing. Let $\mathcal{H}_{0}:\mathcal{C}$ denote the hypothesis that the classifier makes a correct prediction, and $\mathcal{H}_{1}:\neg\mathcal{C}$ the hypothesis that it makes an incorrect one. Selective classification then reduces to deciding, for each input, whether to accept $\mathcal{H}_{0}$ or reject it in favor of $\mathcal{H}_{1}$, i.e., a binary decision problem. This perspective is a natural fit for a foundational result from statistics: the Neyman–Pearson (NP) lemma [14, 15], which characterizes the optimal decision rule between two competing hypotheses.
Lemma 1 (name=Neyman–Pearson [14, 15]).
Let $Z\in\mathbb{R}^{d}$ be a random variable, and consider the hypotheses:
@@
Applied to selective classification, Lemma 1 implies that the optimal selection score takes the form of a likelihood ratio:
@@
Corollary 1 follows directly from the lemma, since monotonic transformations (e.g., logarithmic or affine maps) preserve the ordering of scores and hence do not alter the resulting acceptance region. We define a score function $s({\bm{x}})$ to be Neyman–Pearson optimal if it is a monotonic transformation of the likelihood ratio $p_{c}({\bm{x}})/p_{w}({\bm{x}})$. In practice, the true likelihood ratio is not accessible, so we approximate it or construct a monotonic proxy that captures the posterior odds of a correct versus incorrect prediction for a given input. In what follows, we first reinterpret existing selection scores from the literature as implicit approximations to this likelihood ratio and show conditions under which they are NP optimal. We then introduce two new distance-based selection functions directly inspired by the NP framework. Finally, we propose a hybrid score that linearly combines multiple selectors which we find performs well in practice.
@@ 3.1 Logit-Based Scores as Approximations to Likelihood Ratios
We begin by revisiting two popular confidence scores in classification, Maximum Softmax Probability (MSP) [2] and Raw Logits (RLog) [3], and show that both can be interpreted as approximations to likelihood ratio tests in the sense of the NP lemma. This provides a theoretical justification for their empirical success as selection functions in selective classification.
Let $l^{(1)}\geq l^{(2)}\geq\cdots\geq l^{(K)}$ denote the logits output by a classifier for a sample ${\bm{x}}$, sorted in descending order. Define the corresponding softmax probabilities $d^{(i)}=\mathrm{softmax}(l^{(i)})$. The MSP score is given by $s_{\mathrm{MSP}}({\bm{x}})=d^{(1)}$, while the RLog score is defined as $s_{\mathrm{RLog}}({\bm{x}})=l^{(1)}-l^{(2)}$. MSP has become a standard baseline for OOD detection and selective classification [2, 1], and RLog has been recently proposed as a strong score for selective classification [3].
@@
Assume the classifier is calibrated in the sense that $P(Y=k\mid p_{\theta}(y=k\mid{\bm{x}})=\hat{p})=\hat{p}$ [16]. Then MSP is Neyman–Pearson optimal for selective classification. Moreover, under the additional assumption that the softmax distribution is concentrated on the top two classes, i.e., $L:=\sum_{i\geq 3}d^{(i)}\ll d^{(2)}$, the RLog score is also Neyman–Pearson optimal.
The proof provided in Appendix A shows that under these assumptions, both MSP and RLog are monotonic transformations of the likelihood ratio $p_{c}/p_{w}$, and therefore is NP optimal by Corollary 1. Of course, these assumptions are not always satisfied in practice. Prior work [16] has shown that modern neural classifiers tend to be poorly calibrated, and has proposed post-hoc calibration methods such as temperature scaling. Notably, RLog has been shown to be invariant to temperature scaling [3], making it robust to miscalibration and a compelling choice in practice. This aligns with our empirical findings in Sec. 5, where RLog generally outperforms MSP (which corresponds to temperature scaling with $T=1$). While the effect of calibration is an important factor in logit-based selective classification methods [17, 18], it lies beyond the scope of this work.
@@ 3.2 Neyman–Pearson Optimal Distance Scores
The logit-based scores discussed in the previous section rely on classifier logit calibration, a condition often violated in practice [16]. To avoid this dependency, we consider distance-based methods that make alternative assumptions independent of calibration. As we show below, these methods approximate the likelihood ratio $p_{c}/p_{w}$ by leveraging spatial relationships in feature space.
Two distance methods widely used in OOD detection are the Mahalanobis distance (MDS) [5] and $k$-Nearest Neighbors (KNN) [6]. Both rely on computing distances between a test sample and training features. Specifically, MDS is defined as $s_{\text{MDS}}({\bm{x}})=\max_{i}-(\phi({\bm{x}})-\mu_{i})^{\top}\Sigma^{-1}(\phi({\bm{x}})-\mu_{i})$, where $\phi({\bm{x}})$ denotes the extracted feature of ${\bm{x}}$, typically from the penultimate or final layer of a trained deep network, $\mu_{i}$ is the empirical mean feature of class $i$, and $\Sigma$ is a shared covariance matrix. In contrast, KNN scores inputs by the negative distance to the $k$-th nearest training feature vector. We introduce $\Delta$-MDS and $\Delta$-KNN, which are modified versions of these scores that explicitly incorporate insights from the NP lemma by estimating separate distributions for correctly and incorrectly classified training samples. Figure 1 gives an overview of our approach, and we provide pseudocode for our proposed methods in Appendix C.
@@
The proof is provided in Appendix A, which shows that $s_{\Delta\text{-MDS}}$ is a monotonic transformation of the likelihood ratio $p_{c}/p_{w}$, and thus is NP optimal as per Corollary 1. The Gaussian assumption on feature representations is supported both empirically and theoretically via connections between Gaussian Discriminant Analysis and softmax classifiers [5], making $\Delta$-MDS well-suited for modern deep classifiers trained on standard supervised learning objectives.
@@
The proof is provided in Appendix A. As in previous cases, it relies on showing that $s_{\Delta\text{-KNN}}$ is a monotonic transformation of the likelihood ratio $p_{c}/p_{w}$. Importantly, this result does not require parametric assumptions on the form of $p_{c}$ or $p_{w}$, unlike $\Delta$-MDS. However, it does depend on asymptotic properties of the $k$-nearest neighbor density estimator, and the required conditions on $k$, $N_{c}$, and $N_{w}$ may be difficult to satisfy in finite-sample settings. As such, both methods have their tradeoffs in terms of modeling assumptions.
In practice, we replace the single $k$-th neighbor distance with the average log-distance to the top $k$ neighbors. Specifically, we use: $D_{\text{KNN}}({\bm{x}};A_{c})=-\frac{1}{k}\sum_{i=1}^{k}\log[u_{i}(\phi({\bm{x}}))]$ and $D_{\text{KNN}}({\bm{x}};A_{w})=-\frac{1}{k}\sum_{i=1}^{k}\log[v_{i}(\phi({\bm{x}}))]$. We find that this smoother version improves empirical performance, as shown in our ablation studies in Sec. 5.3. While this modification deviates from the form in Theorem 3, we include a discussion in Appendix B suggesting NP optimality holds for the averaged log-distance formulation under standard assumptions.
@@ 3.3 Linear Combinations Of Distance and Logit-based Scores
The selector scores we have introduced rely on different modeling assumptions and exhibit complementary strengths. Logit-based scores utilize the classifier’s learned boundaries, while distance-based methods depend on geometric structures in feature space defined by training samples. This observation motivates a natural question: can we combine these scores to leverage their respective advantages? We answer in the affirmative by proposing a simple yet effective solution: linearly combining selector scores. Intuitively, this allows each score to compensate for the limitations of the other. The following lemma formalizes the NP optimality of such a linear combination:
@@
The proof provided in Appendix A follows by expressing $t({\bm{x}})$ as a log-product of likelihood ratios. Thus, $t({\bm{x}})$ remains NP optimal under the assumption that the density for each hypothesis takes the form of a multiplicative (or “tilted”) product: $p_{c}^{(1)}(p_{c}^{(2)})^{\lambda}/Z_{c}$ and similarly for $p_{w}$, where $Z_{c}$ is a normalization constant. In practice, we find that combining a distance-based score (e.g., $\Delta$-MDS) with a logit-based score (e.g., RLog) leads to the best performance. We refer to such combinations by concatenating their names, e.g., $\Delta$-MDS-RLog. While the mixing coefficient $\lambda$ can be determined via grid search, we observe that balancing the magnitudes of $s_{1}({\bm{x}})$ and $s_{2}({\bm{x}})$ yields the best results without requiring extensive tuning.
@@ 4 Related Works
The study of classification with a reject option is long-standing. Early foundational works considered cost-based rejection models [10], as well as in classical machine learning paradigms such as support vector machines [19, 20] and nearest neighbor estimators [21]. In the context of deep learning, LeCun et al. [22] was among the first to examine rejection based on the top logits’ activations in image classification. A modern treatment of selective classification was later developed [23, 1], introducing the key concepts of classifier–selector pairs and the risk–coverage trade-off, which underpin our framework. Geifman and El-Yaniv [1] further proposed a method for achieving guaranteed risk bounds and evaluated selection strategies based on model confidence scores, such as softmax response (equivalent to MSP) and Monte Carlo dropout [4]. These approaches share the underlying assumption that confidence estimates are well-calibrated indicators of predictive correctness.
@@
Subsequent works have extended this direction by studying popular logit and distance-based scores, many originally developed for OOD detection. Notable examples include MSP [2], MaxLogit [24], Energy [25], MDS [5], and KNN [6]. A critical limitation of logit-based methods is their reliance on classifier calibration. Although calibration is not the focus of our work, several studies have examined its impact on selective classification performance [17, 26]. Related to this, conformal prediction provides an alternative framework for producing calibrated confidence or prediction sets with formal coverage guarantees [27, 28, 29, 30]. While such methods could in principle be adapted for selective classification, they differ fundamentally from our goal of designing scoring functions optimized for selective risk.
Beyond post-hoc methods, training-time approaches have been proposed where models are explicitly designed with a reject option. For example, SelectiveNet [7] incorporates a rejection head and jointly trains the model to both classify and reject. Similarly, Deep Gamblers [8] and Self-Adaptive Training [9] reformulate $K$-way classification into $K+1$ classes by introducing an explicit reject class. These methods differ from ours in that they require architectural modifications and joint training. In contrast, we focus on post-hoc methods that can be applied to any pretrained classifier without retraining. This enhances applicability in practice, where most deployed models do not include rejection mechanisms by design.
@@
Most closely related to our work is Liang et al. [3], who study selective classification under general distribution shifts, which include both semantic and covariate shifts. They also introduce the Raw Logit (RLog) score, which we analyze in this work. Our work differs in several key ways: 1) we focus specifically on covariate shifts, which we argue is more relevant in modern settings where large and variable label sets (e.g., from vision-language models) mitigate label drift; 2) we introduce a unified theoretical framework grounded in the Neyman–Pearson lemma, from which we derive new selector scores with formal optimality guarantees; and 3) we evaluate our methods on a broader class of models, including VLMs, whereas Liang et al. [3] focus exclusively on standard supervised learning paradigms.

### v2

@@ 1 Introduction
A wide range of methods have been proposed to determine whether a model should accept or reject an input. Common post-hoc approaches rely on heuristic confidence estimates, such as the maximum softmax probability (Geifman and El-Yaniv, 2017; Hendrycks and Gimpel, 2017), logit margins (Liang et al., 2024), or Monte Carlo dropout (Geifman and El-Yaniv, 2017; Gal and Ghahramani, 2016). Other techniques assess a sample’s proximity to the training distribution (Lee et al., 2018; Sun et al., 2022), under the assumption that samples farther from the data manifold are more likely to be misclassified. A separate line of work trains models with explicit abstention mechanisms, such as a rejection logit or head (Geifman and El-Yaniv, 2019; Liu et al., 2019; Huang et al., 2020). In this work we focus on post-hoc methods, which are model-agnostic and do not require specialized training.
Despite the rich literature, two important gaps remain. First, while foundational results (such as Chow (1970); Geifman and El-Yaniv (2017)) provide theoretical underpinnings for selective classification, there is a lack of general, principled guidance for designing effective selector functions in the context of modern deep networks. Second, most evaluations are conducted in the i.i.d. setting, where test data is assumed to follow the training distribution. Few works have begun exploring selective classification under distribution shifts (Xia and Bouganis, 2022; Narasimhan et al., 2024), but focus on the common semantic shifts, neglecting the covariate shift setting that is becoming increasingly relevant.
To address these challenges, we propose a new perspective rooted in the Neyman–Pearson lemma, a classical result from statistics that defines the optimal hypothesis test in terms of a likelihood ratio. We show that existing selectors can be interpreted as approximations to this test, and we use this insight to derive two new selectors, $\Delta$-MDS and $\Delta$-KNN, as well as a simple linear combination strategy. We evaluate our methods on a comprehensive suite of vision and language benchmarks under covariate shift, where the input distribution changes while the label space remains fixed. We focus on covariate shift for two key reasons: first, it is underexplored relative to semantic shifts (Heng and Soh, 2025) (which is well studied in the context of out-of-distribution detection (Hendrycks and Gimpel, 2017; Ming et al., 2022; Lee et al., 2018; Heng et al., 2024)); second, it is increasingly relevant in modern applications such as vision-language models (VLMs), where the label set is large and variable, rendering most practical shifts in deployment covariate in nature. Our results demonstrate that the proposed selectors outperform existing baselines and provide robust performance across distribution shifts, including on powerful VLMs like CLIP.
@@
    We conduct a thorough evaluation under distribution shifts, both covariate and semantic, across vision and language tasks and demonstrate superior performance across both VLMs and traditional supervised models.
@@ 2 Background
where $s:\mathcal{X}\rightarrow\mathbb{R}$ is a confidence scoring function (often adapted from OOD detection methods; see below), and $\gamma$ is a tunable threshold. The performance of a selective classifier is typically evaluated using two metrics:
@@
where $\ell(f({\bm{x}}),y)$ is the $0/1$ loss (Geifman and El-Yaniv, 2017). Selective classification aims to optimize the tradeoff between selective risk and coverage by ideally reducing risk while maintaining high coverage. Improvements can stem from enhancing the base classifier $f$, or from refining the selector $g$ to better identify error-prone inputs. In this work, we fix $f$ to be a strong pretrained model and focus on designing more effective selector functions $g$.
@@ 3 Selective Classification via the Neyman–Pearson Lemma
We begin by framing selective classification within the paradigm of hypothesis testing. Let $\mathcal{H}_{0}:\mathcal{C}$ denote the hypothesis that the classifier makes a correct prediction, and $\mathcal{H}_{1}:\neg\mathcal{C}$ that it makes an incorrect one. Selective classification then reduces to deciding, for each input, whether to accept $\mathcal{H}_{0}$ or reject in favor of $\mathcal{H}_{1}$, i.e., a binary decision problem. This perspective is a natural fit for a foundational result from statistics: the Neyman–Pearson (NP) lemma (Neyman and Pearson, 1933; Lehmann et al., 1986), which characterizes the optimal decision rule between two competing hypotheses.
Lemma 1 (name=Neyman–Pearson (Neyman and Pearson, 1933; Lehmann et al., 1986)).
Let $Z\in\mathbb{R}^{d}$ be a random variable, and consider the hypotheses::
@@
Applied to selective classification, Lemma 1 implies the optimal selection score is a likelihood ratio:
@@
Corollary 1 follows directly from the lemma, since monotonic transformations (e.g., logarithmic or affine maps) preserve the ordering of scores and hence do not alter the resulting acceptance region. We define a score function $s({\bm{x}})$ to be Neyman–Pearson optimal if it is a monotonic transformation of the likelihood ratio $p_{c}({\bm{x}})/p_{w}({\bm{x}})$. In practice, the true likelihood ratio is not accessible, so we approximate it or construct a monotonic proxy that captures the posterior odds of a correct versus incorrect prediction for a given input.
Note that $p_{c}({\bm{x}})$ and $p_{w}({\bm{x}})$ are general and naturally accounts for distribution shifts; $p_{c}({\bm{x}})$ ($p_{w}({\bm{x}})$) includes all samples that the classifier classifies correctly (wrongly), regardless of whether they are ID or a distribution shift. Therefore, this simplifies our framework compared to prior works that consider ID and OOD distributions separately (Xia and Bouganis, 2022; Narasimhan et al., 2024).
In what follows, we first reinterpret existing selection scores from the literature as implicit approximations to this likelihood ratio and show conditions under which they are NP optimal. We then introduce two new distance-based selection functions inspired by the NP framework. Finally, we propose a hybrid score that linearly combines multiple selectors which we find performs well in practice. Throughout, the assumptions used in our theoretical results are meant to clarify the connection to NP optimality and the structure of an optimal selector, rather than to prescribe conditions that must hold in practice. Introduction and brief details of the scores discussed in this section are provided in Appendix A.
@@ 3.1 Logit-Based Scores as Approximations to Likelihood Ratios
We consider two popular confidence scores, Maximum Softmax Probability (MSP) (Hendrycks and Gimpel, 2017) and Raw Logits (RLog) (Liang et al., 2024), and interpret them as approximations to likelihood ratio tests in the NP sense. This provides theoretical justification for their empirical success as selector scores in selective classification.
Let $l^{(1)}\geq l^{(2)}\geq\cdots\geq l^{(K)}$ denote the logits output by a classifier for a sample $\bm{x}$}, sorted in descending order. Define the corresponding softmax probabilities $d^{(i)}=\mathrm{softmax}(l^{(i)})$. The MSP score is given by $s_{\mathrm{MSP}}({\bm{x}})=d^{(1)}$, while the RLog score is defined as $s_{\mathrm{RLog}}({\bm{x}})=l^{(1)}-l^{(2)}$. MSP has become a standard baseline for OOD detection and selective classification (Hendrycks and Gimpel, 2017; Geifman and El-Yaniv, 2017), and RLog has been recently proposed as a strong score for selective classification (Liang et al., 2024).
@@
Assume the classifier is calibrated in the sense that $P(Y=k\mid p_{\theta}(y=k\mid{\bm{x}})=\hat{p})=\hat{p}$ (Guo et al., 2017). Then MSP is Neyman–Pearson optimal for selective classification. Moreover, under the additional assumption that the softmax distribution is concentrated on the top two classes, i.e., $L:=\sum_{i\geq 3}d^{(i)}\ll d^{(2)}$, the RLog score is also Neyman–Pearson optimal.
The proof provided in Appendix B shows that under these assumptions, both MSP and RLog are monotonic transformations of the likelihood ratio $p_{c}/p_{w}$, and therefore is NP optimal by Corollary 1. Of course, these assumptions are not always satisfied in practice. Prior work (Guo et al., 2017) has shown that modern neural classifiers tend to be poorly calibrated, and has proposed post-hoc calibration methods such as temperature scaling. Notably, RLog has been shown to be invariant to temperature scaling (Liang et al., 2024), making it robust to miscalibration and a compelling choice in practice. This aligns with our empirical findings in Sec. 5, where RLog generally outperforms MSP (which corresponds to temperature scaling with $T=1$). While the effect of calibration is an important factor in logit-based methods (Cattelan and Silva, 2023; Fisch et al., 2022), it lies beyond the scope of this work.
@@ 3.2 Neyman–Pearson Optimal Distance Scores
The logit-based scores discussed in the previous section rely on classifier logit calibration, a condition often violated in practice (Guo et al., 2017). To avoid this dependency, we consider distance-based methods that make alternative assumptions independent of calibration. As we show below, these methods approximate the likelihood ratio $p_{c}/p_{w}$ by leveraging spatial relationships in feature space.
Two distance methods widely used in OOD detection are the Mahalanobis distance (MDS) (Lee et al., 2018) and $k$-Nearest Neighbors (KNN) (Sun et al., 2022). Both rely on computing distances between a test sample and training features (see Appendix A for details). Briefly, MDS is defined as $s_{\text{MDS}}({\bm{x}})=\max_{i}-(\phi({\bm{x}})-\mu_{i})^{\top}\Sigma^{-1}(\phi({\bm{x}})-\mu_{i})$, where $\phi({\bm{x}})$ denotes the extracted feature of $\bm{x}$}, typically from the penultimate or final layer of a trained deep network, $\mu_{i}$ is the empirical mean feature of class $i$, and $\Sigma$ is a shared covariance matrix. In contrast, KNN scores inputs by the negative distance to the $k$-th nearest training feature vector. We introduce $\Delta$-MDS and $\Delta$-KNN, which are modified versions of these scores that explicitly incorporate insights from the NP lemma by estimating separate distributions for correctly and incorrectly classified training samples. Figure 1 gives an overview of our approach, and we provide pseudocode for our proposed methods in Appendix D.
@@
The proof is provided in Appendix B, which shows that $s_{\Delta\text{-MDS}}$ is a monotonic transformation of the likelihood ratio $p_{c}/p_{w}$, and thus is NP optimal as per Corollary 1. The Gaussian assumption on feature representations is supported both empirically and theoretically via connections between Gaussian Discriminant Analysis and softmax classifiers (Lee et al., 2018), making $\Delta$-MDS well-suited for modern deep classifiers trained on standard supervised learning objectives.
@@
The proof is provided in Appendix B. As in previous cases, it relies on showing that $s_{\Delta\text{-KNN}}$ is a monotonic transformation of the likelihood ratio $p_{c}/p_{w}$. Importantly, this result does not require parametric assumptions on the form of $p_{c}$ or $p_{w}$, unlike $\Delta$-MDS. However, it does depend on asymptotic properties of the $k$-nearest neighbor density estimator, and the required conditions on $k$, $N_{c}$, and $N_{w}$ may be difficult to satisfy in finite-sample settings. As such, both methods have their tradeoffs in terms of modeling assumptions.
In practice, we replace the single $k$-th neighbor distance with the average log-distance to the top $k$ neighbors. Specifically, we use: $D_{\text{KNN}}({\bm{x}};A_{c})=-\frac{1}{k}\sum_{i=1}^{k}\log[u_{i}(\phi({\bm{x}}))]$ and $D_{\text{KNN}}({\bm{x}};A_{w})=-\frac{1}{k}\sum_{i=1}^{k}\log[v_{i}(\phi({\bm{x}}))]$. We find that this smoother version improves empirical performance, as shown in our ablation studies in Sec. 5.3. While this modification deviates from the form in Theorem 3, we include a discussion in Appendix C suggesting NP optimality holds for the averaged log-distance formulation under standard assumptions.
@@ 3.3 Linear Combinations Of Distance and Logit-based Scores
The selector scores we discussed rely on different modeling assumptions and exhibit complementary strengths, as discussed in Wang et al. (2022). Logit-based scores utilize the classifier’s learned boundaries, while distance-based methods depend on geometric structures in feature space defined by training samples. We are thus motivated to leverage their respective advantages by proposing a simple yet effective solution: linearly combining selector scores. Intuitively, this allows each score to compensate for the limitations of the other. The following lemma formalizes the NP optimality of such a linear combination:
@@
The proof provided in Appendix B follows by expressing $t({\bm{x}})$ as a log-product of likelihood ratios. Thus, $t({\bm{x}})$ remains NP optimal under the assumption that the density for each hypothesis takes the form of a multiplicative (or “tilted”) product: $p_{c}^{(1)}(p_{c}^{(2)})^{\lambda}/Z_{c}$ and similarly for $p_{w}$, where $Z_{c}$ is a normalization constant. In practice, we find that combining a distance-based score (e.g., $\Delta$-MDS) with a logit-based score (e.g., RLog) leads to the best performance. We refer to such combinations by concatenating their names, e.g., $\Delta$-MDS-RLog. We discuss fitting parameters like $\lambda$ in Sec. 5.
@@ 4 Related Works
The study of classification with a reject option has a long history, beginning with cost-based formulations (Chow, 1970) and extensions to classical models like SVMs (Fumera and Roli, 2002; Bartlett and Wegkamp, 2008) and nearest neighbors (Hellman, 1970). In deep learning, LeCun et al. (1989) explored rejection via top logit activations. Later, the risk–coverage and classifier–selector frameworks were formalized (El-Yaniv and others, 2010; Geifman and El-Yaniv, 2017), with methods like MSP and Monte Carlo dropout proposed to provide confidence-based selection (Gal and Ghahramani, 2016).
@@
Subsequent works have extended this direction by studying popular logit and distance-based scores, many originally developed for OOD detection. Examples include MSP (Hendrycks and Gimpel, 2017), MaxLogit (Hendrycks et al., 2019), Energy (Liu et al., 2020), MDS (Lee et al., 2018), and KNN (Sun et al., 2022). A limitation of logit-based methods is their reliance on classifier calibration. Although calibration is not the focus of our work, several studies have examined its impact on selective classification performance (Cattelan and Silva, 2023; Galil et al., 2023). An alternative line of research uses conformal prediction to construct calibrated prediction sets with formal guarantees (Vovk et al., 2005; Angelopoulos and Bates, 2021; Bates et al., 2021; Angelopoulos et al., 2024). While such methods could be adapted for selective classification, they differ fundamentally from our goal of designing scoring functions optimized for selective risk. Jiang et al. (2018) proposes a trust score by comparing the model’s prediction with class-conditional KNN distances to estimate if a sample will be correctly classified, but it was shown to be ineffective on high-dimensional images.
Some methods incorporate rejection directly into training. For instance, SelectiveNet (Geifman and El-Yaniv, 2019) adds a dedicated rejection head, while Deep Gamblers (Liu et al., 2019) and Self-Adaptive Training (Huang et al., 2020) introduce a reject class and train the model to abstain. These methods require architectural modifications and joint training. In contrast, our work focuses on post-hoc methods that can be applied to pretrained classifiers without retraining.
Related to our work are approaches that combine selective classification and OOD detection, termed SCOD (Xia and Bouganis, 2022; Narasimhan et al., 2024). Such methods consider the ID classification and OOD distributions separately and seek to combine them into a single score function. These approaches are typically designed for semantic shifts and need adaptation for covariate shift. Our formulation avoids this by representing all distribution shifts through the general pair $(p_{c},p_{w})$, which does not require distinguishing between shift types.
@@
Finally, most closely related to our work is Liang et al. (2024), who study selective classification under both semantic and covariate shifts and introduce the Raw Logit (RLog) score. Our work differs in several ways: 1) we focus on covariate shifts, which we argue is more relevant in modern settings where large and variable label sets (e.g., from vision-language models) mitigate label drift; 2) we introduce a unified theoretical framework grounded in the Neyman–Pearson lemma, from which we derive new selector scores with formal optimality guarantees; and 3) we evaluate our methods on a broader class of models, including VLMs, whereas Liang et al. (2024) focus exclusively on standard supervised learning paradigms.